OpenStudy (anonymous):

Let f(t)=t^2+t. What is the average rate of change in f(t) between t=2 and t=3? Show all work.

jimthompson5910 (jim_thompson5910):

What is f(2) equal to?

OpenStudy (anonymous):

It doesn't say. I'm guessing it just uses the equation f(b)-f(a)/b-a?

jimthompson5910 (jim_thompson5910):

you have to determine that bit of info

jimthompson5910 (jim_thompson5910):

so plug t = 2 into the function and evaluate

OpenStudy (anonymous):

(3^2 + 3) - (2^2 +2) / 3-2 = (9+3) - (4+3) / 3-2 = (12 - 7) / 1 = 5/1 = 5 Does that look right?

jimthompson5910 (jim_thompson5910):

you somehow went from 2^2 +2 to 4 +3

jimthompson5910 (jim_thompson5910):

when it should be 2^2 +2 ----> 4 + 2

jimthompson5910 (jim_thompson5910):

otherwise, it looks good

OpenStudy (anonymous):

(3^2 + 3) - (2^2 +2) / 3-2 = (9+3) - (4+2) / 3-2 = (12 - 6) / 1 = 6/1 = 5 Woops, I guess I was thinking of the t=3 or something and got momentarily confused >.<

jimthompson5910 (jim_thompson5910):

I'm assuming you meant to say =6 on the last line, but you have the answer the average rate of change is 6

OpenStudy (anonymous):

6/1 = 6

OpenStudy (anonymous):

Yeah. I guess I'm just bad and typing math out or something, haha. Thanks!

jimthompson5910 (jim_thompson5910):

no worries and you're welcome